### 25 Dec 2023 v1

# Privacy-Preserving Neural Graph Databases

Abstract.
In the era of big data and rapidly evolving information systems, efficient and accurate data retrieval has become increasingly crucial. Neural graph databases (NGDBs) have emerged as a powerful paradigm that combines the strengths of graph databases (graph DBs) and neural networks to enable efficient storage, retrieval, and analysis of graph-structured data. The usage of neural embedding storage and complex neural logical query answering provides NGDBs with generalization ability. When the graph is incomplete, by extracting latent patterns and representations, neural graph databases can fill gaps in the graph structure, revealing hidden relationships and enabling accurate query answering. Nevertheless, this capability comes with inherent trade-offs, as it introduces additional privacy risks to the database. Malicious attackers can infer more sensitive information in the database using well-designed combinatorial queries, such as by comparing the answer sets of where Turing Award winners born before 1950 and after 1940 lived, the living places of Turing Award winner Hinton are probably exposed, although the living places may have been deleted in the training due to the privacy concerns. In this work, inspired by the privacy protection in graph embeddings, we propose a privacy-preserving neural graph database (P-NGDB) to alleviate the risks of privacy leakage in NGDBs. We introduce adversarial training techniques in the training stage to force the NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring sensitive information through combinations of multiple innocuous queries. Extensive experiment results on three datasets show that P-NGDB can effectively protect private information in the graph database while delivering high-quality public answers responses to queries.
privacy preserving, knowledge graph, complex query answering
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Privacy protectionsccs: Information systems Data mining
1. Introduction
Graph databases (graph DBs) play a crucial role in storing, organizing, and retrieving structured relational information and are widely applied in downstream applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. Unfortunately, traditional graph databases often suffer from the limitation of graph incompleteness, which is a prevalent issue in real-world knowledge graphs, like Freebase (Mintz et al., 2009; Bollacker et al., 2008). This incompleteness leads to the exclusion of relevant results, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013). To solve the drawbacks, extending the concept of Graph DBs, neural graph databases (NGDBs) combine the flexibility of graph data models with the computational capabilities of neural networks, enabling efficient representation, storage, and analysis of interconnected data (Ren et al., 2023; Hamilton et al., 2018).
Figure 1. Structure of Neural Graph Databases (NGDBs). NGDBs can retrieve answers given complex logical queries on incomplete graphs. However, the privacy problem arises due to the storage of graph information in embedding space and the presence of a query engine that facilitates complex logical query answering. To illustrate the issue, consider an example where an attacker attempts to infer private information about Hinton’s living place in the NGDBs. Direct querying privacy can be easily detected by privacy risk detection, however, attackers can leverage the combination of multiple queries to retrieve desired privacy.
NGDBs provide unified storage for diverse entities in an embedding space. By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020). As shown in Figure 1, logical expressions are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. The logical forms of $q$ include terms like $Win(V,Turing\ Award)$ and logical operators like conjunction $\wedge$, representing that find where the Turing Award winner who was born in 1960 lived. Query encoding methods are commonly used in Complex Query Answering (CQA). These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query embedding and entity embedding. For example, GQE (Hamilton et al., 2018) and query2box (Ren et al., 2020) encode queries as vector and hyper-rectangle, respectively.
While NGDBs have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge graphs and enriches the retrieval information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. As shown in Figure 1, the living place information of an individual, in this case, Hinton as an example, is considered private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place. For example, by querying NGDBs with where Turing Award winners born before 1950 and after 1940 lived, where LeCun’s collaborator lived, etc. The intersection of these queries still presents a significant likelihood of exposing the living place of Turing Award winner Hinton.
Graph privacy has consistently been a significant concern in the field of graph research. Multiple studies demonstrate the vulnerability of graph neural networks to various privacy leakage issues. For instance, graph embeddings are vulnerable to attribute inference attacks (Duddu et al., 2020; Gong and Liu, 2018; Zhang et al., 2022), link prediction attacks (He et al., 2021b; Wu et al., 2022), etc. and various protection methods are proposed (Hu and Song, 2023; Hu et al., 2022; Wang et al., 2021). However, these works focus on graph representation learning and have the assumption that the graph embeddings can be reached by the attackers which are not typical scenarios in NGDBs. NGDBs often provide CQA service and do not publish the learned embeddings, and attackers only can infer privacy using thoughtfully designed queries. Furthermore, these works solely study node classification and link prediction problems, leaving the query answering privacy in NGDBs unexplored.
In this paper, we expose the potential privacy risks of NGDBs with formal definition and evaluation. We introduce privacy protection objectives to the existing NGDBs, categorizing the answers of knowledge graph complex queries into private and public domains. In order to safeguard the private information within the knowledge graph, NGDBs should strive to maintain high-quality retrieval of non-private answers for queries, while intentionally obfuscating the private-threatening answers to innocuous yet private queries. As shown in figure 1, the answer sets of queries proposed by privacy attackers are obfuscated, and attackers will face greater challenges in predicting private information, thereby effectively safeguarding the privacy of NGDBs. Meanwhile, we also create the corresponding benchmark on three datasets (Freebase (Bollacker et al., 2008), YAGO (Suchanek et al., 2007), and DBpedia (Bizer et al., 2009)) to evaluate the performance of CQA ability on public query answers and the protection of private information.
To alleviate the privacy leakage problem in NGDBs, we propose Privacy-preserving Neural Graph Databases (P-NGDBs) as a solution. P-NGDBs divide the information in graph databases into private and public parts and categorize complex queries’ answers to various privacy risks based on whether or not they involve sensitive information. P-NGDBs can provide answers with different levels of precision in response to queries with varying privacy risks. We introduce adversarial techniques in the training stage of P-NGDBs to generate indistinguishable answers when queried with private information, enhancing the difficulty of inferring privacy through complex private queries. We summarize our major contributions as follows:
  • To the best of our knowledge, our work represents the pioneering effort in investigating privacy leakage issues of complex queries in NGDBs and providing formal definitions for privacy protection in this domain.
  • Based on three public datasets, we propose a benchmark for the simultaneous evaluation of CQA ability and privacy protection for NGDBs.
  • We introduce P-NGDB, a privacy protection framework for safeguarding privacy in NGDBs. Extensive experiments conducted on three datasets demonstrate its ability to effectively preserve high information retrieval performance while ensuring privacy protection.
The rest of the paper is organized as follows. Section 2 systematically reviews the related work. Section 3 introduces the preliminary and definition of the privacy problem in NGDBs. Section 4 and 5 introduce the innocuous yet private queries and the framework of P-NGDBs in detail, respectively. Section 6 evaluates the performance of P-NGDBs on the benchmark based on three real-world datasets. Finally, we conclude our work in Section 7.
2. Related Work
In this section, we summarize the related work. Our work is related to complex query answering (CQA) and graph privacy.
2.1. Complex Query Answering
One essential function of neural graph databases is to answer complex structured logical queries according to the data stored in the NGDBs, namely complex query answering (CQA). Query embedding is one of the prevalent methods for complex query answering because of its effectiveness and efficiency. These methods for query embedding utilize diverse structures to represent complex logical queries and effectively model queries across various scopes. Some methods encode queries into various geometric representations: The GQE (Hamilton et al., 2018) model encodes queries to vectors in embedding space, Query2box (Ren et al., 2020) encodes queries to hyper-rectangle embeddings, HypeE (Liu et al., 2021) encodes queries as hyperbolic embeddings to further support disjunction logical operators; Query2Particle (Bai et al., 2022), ConE (Zhang et al., 2021b) supports negation operators using multiple vector embeddings and cone embeddings, respectively. NRN (Bai et al., 2023a) is proposed to encode numerical values. Besides, some methods try to use probabilistic structures to encode logic queries: BetaE (Ren and Leskovec, 2020), GammaE (Yang et al., 2022), PERM (Choudhary et al., 2021) propose to use Beta, Gamma, and Gaussian distributions to encode logic knowledge graph queries, respectively. Meanwhile, Neural structures are utilized to encode complex queries: BiQE (Kotnis et al., 2021), SQE (Bai et al., 2023b), QE-GNN (Zhu et al., 2022) are proposed to use transformer, sequential encoders, and message-passing graph neural networks to encode queries, respectively. Query decomposition (Arakelyan et al., 2020) proposes to decompose complex queries and LMPNN (Wang et al., 2022) proposes to use one-hop message passing on query graphs to conduct CQA.
While complex query answering has been widely studied in past works, the privacy problem in NGDBs is overlooked. With the development of the representation ability in NGDBs, the issue of privacy leakage has become increasingly critical. There are some works showing that graph representation learning is vulnerable to various privacy attacks (Hu et al., 2022). (Salem et al., 2018; Olatunji et al., 2021) show that membership inference attacks can be applied to identify the existence of training examples. Model extraction attacks (Tramèr et al., 2016) and link stealing attacks (He et al., 2021b; Zhang et al., 2021a) try to infer information about the graph representation model and original graph link, respectively. However, these works assume that attackers have complete access to the graph representation models. In NGDBs, however, sensitive information can be leaked during the query stage, presenting a different privacy challenge. Our research explores the issue of privacy leakage in NGDBs and presents a benchmark for comprehensive evaluation.
2.2. Graph Privacy
Various privacy protection methods have been proposed to preserve privacy in graph representation models. Anonymization techniques (Liu and Terzi, 2008; Rossi et al., 2015; Hoang et al., 2020; Zheleva and Getoor, 2007) are applied in graphs to reduce the probability of individual and link privacy leakage. Graph summarization (Hay et al., 2008) aims to publish a set of anonymous graphs with private information removed. Learning-based methods are proposed with the development of graph representation learning, (Wang et al., 2021; Liao et al., 2020; Li et al., 2020) regard graph representation learning as the combination of two sub-tasks: primary objective learning and privacy preservation, and use adversarial training to remove the sensitive information while maintaining performance in the original tasks. Meanwhile, some methods (Liu et al., 2022; Oh et al., 2022; Hu and Song, 2023) disentangle the sensitive information from the primary learning objectives. Additionally, differential privacy (Kasiviswanathan et al., 2013; Day et al., 2016; Shen and Yu, 2013) introduces noise into representation models and provides privacy guarantees for the individual privacy of the datasets (Daigavane et al., 2021). Though noise introduced by differential privacy protects models from privacy leakage, it can also impact the performance of the original tasks significantly. Federated learning utilizes differential privacy and encryption to prevent the transmission of participants’ raw data and is widely applied in distributed graph representation learning (Peng et al., 2021; He et al., 2021a). However, federated learning cannot be applied to protect intrinsic private information in the representation models.
While there are various graph privacy preservation methods, they only focus on simple node-level or edge-level privacy protection. However, the privacy risks associated with neural graph databases during complex query answering tasks have not received sufficient research attention. The development of CQA models introduces additional risks of privacy leakage in NGDBs that attackers can infer sensitive information with multiple compositional queries. Our proposed P-NGDB can effectively reduce the privacy leakage risks of malicious queries.
3. Preliminary and Problem Formulation
3.1. Preliminary
We consider the privacy problem in the multi-relational knowledge graph with numerical attributes. We denote a knowledge graph as $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A},\mathcal{N})$, where $\mathcal{V}$ denotes the set of vertices representing entities in the knowledge graph. $\mathcal{R}$ denotes the set of relations where $r\in\mathcal{R}$ is a binary function defined as $r:\mathcal{V}\times\mathcal{V}\rightarrow\{0,1\}$ describing the relation between entities in the knowledge graph, where $r(u,v)=1$ denotes there is a relation $r$ between entities $u$ and $v$ and $r(u,v)=0$ otherwise. $\mathcal{A}$ denotes the set of attributes that can be split into private attributes and public attributes: $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a\in\mathcal{A}$ is a binary function defined as $a:\mathcal{V}\times\mathcal{N}\rightarrow\{0,1\}$ describing the numerical attribute of entities, where $a(u,x)=1$ denotes the entity $u$ has the attribute $a$ of value $x\in\mathcal{N}$ and $a(u,x)=0$ otherwise. Besides, there are $f(x_{1},x_{2}):\mathcal{N}\times\mathcal{N}\rightarrow\{0,1\}$ to denote the relation between numerical values. To prevent privacy leakage, the private attribute of entities $\mathcal{A}_{\text{private}}$ cannot be exposed and should be protected from inferences.
The complex query is the key task of NGDBs and can be defined in existential positive first-order logic form, consisting of various types of logic expressions like existential quantifiers $\exists$, logic conjunctions $\wedge$, and disjunctions $\vee$. In the logical expression, there is a unique variable $V_{?}$ in each logic query that denotes our query target. Variables $V_{1},\cdots,V_{k}$ and numerical variables $X_{1},\cdots,X_{l}$ denote existential quantified entities in a complex query. Besides, there are anchor entities $V_{a}$ and values $X_{a}$ with the given content in a query. The complex query aims to identify the target entity $V_{?}$, such that there are $V_{1},\cdots,V_{k}\in\mathcal{V}$ and $X_{1},\cdots,X_{l}\in\mathcal{N}$ in the knowledge graph that can satisfy the given logical expressions. We denote complex query expression in the disjunctive normal form (DNF) in the following:
(1) $\begin{split}q[V_{?}]&=V_{?}.V_{1},...,V_{k},X_{1},...,X_{l}:c_{1}\lor c_{2}\lor...\lor c_{n}\\ c_{i}&=e_{i,1}\land e_{i,2}\land...\land e_{i,m},\end{split}$
where $e_{i,j}$ is the atomic logic expression, which can be $r(V,V^{\prime})$ denotes relation $r$ between entities $V$ and $V^{\prime}$, or $a(V,X)$ denotes attribute $a$ of entities $V$ with value $X$ and $f(X,X^{\prime})$ denotes the relation between numerical attributes. $c_{i}$ is the conjunction of several atomic logic expressions $e_{i,j}$. $V,V^{\prime},X$ are either anchor entities or values, or existentially quantified variables or numearical variables.
3.2. Problem Formulation
Suppose that there is a graph $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{A},\mathcal{N})$, where part of entities’ attributes are regarded as private information $\mathcal{A}=\mathcal{A}_{\text{private}}\cup\mathcal{A}_{\text{public}}$, where $a_{p}\in\mathcal{A}_{\text{private}}$ denotes the entity $u$ has the sensitive attribute $a_{p}$ with value $x$ if $a_{p}(u,x)=1$. Due to the variation in privacy requirements, an attribute can be handled differently for various entities. NGDBs store the graph in embedding space and can be queried with arbitrary complex logical queries. Due to the generalization ability of NGDBs, attackers can easily infer sensitive attributes $\mathcal{A}_{\text{private}}$ utilizing complex queries. To protect private information, NGDBs should retrieve obfuscated answers when queried with privacy risks while preserving high accuracy in querying public information.
4. Innocuous yet Private Query
As there is sensitive information in the neural graph databases, some specified complex queries can be utilized by attackers to infer privacy, which is defined as Innocuous yet Private (IP) queries. Some answers to these queries can only be inferred under the involvement of private information, which is a risk to privacy and should not be retrieved by the NGDBs.
4.1. Computational Graph
The query $q$ can be parsed into a computational graph as a directed acyclic graph, which consists of various types of directed edges that represent operators over sets of entities and attributes.
  • Projection: Projection has various types. Given a set of entities $S$, a relation between entities $r\in\mathcal{R}$, an attribute type $a$, a set of values $S\subset\mathcal{N}$, projection describes the entities or values can be achieved under specified relation and attribute types. For example, attribute projection is denoted as $P_{a}(S)=\{x\in\mathcal{N}|v\in S,a(v,x)=1\}$ describing the values that can be achieved under the attribute type $a$ by any of the entities in $S$. Furthermore, we have relational projection $P_{r}(S)$, reverse attribute projection $P_{a}^{-1}(S)$ and numerical projection $P_{f}(S)$. We denote all these projections as $P(S)$;
  • Intersection: Given a set of entity or value sets $S_{1},S_{2},...,S_{n}$, this logic operator computes the intersection of those sets as $\cap_{i=1}^{n}S_{i}$;
  • Union: Given a set of entity or value sets $S_{1},S_{2},...,S_{n}$, the operator computes the union of $\cup_{i=1}^{n}S_{i}$
4.2. Privacy Threatening Query Answers
After a complex query is parsed into the computational graph $G$, and $M=G(S)$ denotes the answer set of entities or values retrieved by the computational graph. We divide the $M$ retrieved by the complex queries as private answer set $M_{\text{private}}$ and non-private answer set $M_{\text{public}}$ based on whether the answer has to be inferred under the involvement of private information. As shown in Figure 2, query in (A) retrieves answers from the knowledge graph, Toronto is regarded as privacy-threatening answers because $LiveIn(Hinton,Toronto)\in\mathcal{A}$.
Projection: A projection operator’s output will be regarded as private if the operation is applied to infer private attributes. Assume that $M=P(S)$ and input $S=S_{\text{private}}\cup S_{\text{public}}$. $P(S)=P(S_{\text{private}})\cup P(S_{\text{public}})$. Without loss of generalization, we discuss the formal definition of projection private answer set $M_{\text{private}}$ in two scenarios. For the projection on public inputs:
$M_{\text{private}}=\{m\mid\exists v\in S:m\in M\land P(v,m)\in\mathcal{A}_{\text{private}}\land$
$P(u,m)\neq P(v,m),\forall u\in S/\{v\}\}.$
For the projection on private inputs,
$M_{\text{private}}=P(S_{\text{Private}}).$
As shown in Figure 3(A), all the red nodes within the answer sets can solely be deduced from private attribute links and are classified as query answers posing privacy risks. On the other hand, nodes such as the green-red node, while accessible through private projection, can also be inferred from public components, thereby carrying lower privacy risks.
Figure 2. An example of query demonstrating the retrieved privacy-threatening query answers. Red node denotes privacy risks. (A) The logic knowledge graph query involves privacy information. (B) Example of a complex query in the knowledge graph, Toronto is regarded as a privacy-threatening answer as it has to be inferred by sensitive information.
Figure 3. Example of privacy-threatening answer sets computation in projection, intersection, and union. Green nodes denote non-private answers, red nodes denote privacy-threatening answers, green-red nodes denote different privacy risks in subsets. Red dashed arrows denote privacy projection. The range demarcated by the red dashed lines denotes privacy-threatening answer sets.
Intersection: An intersection operator’s output will be regarded as private if the answer belongs to any of the private answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, where each answer set $M^{i}=M^{i}_{\text{private}}\cup M^{i}_{\text{public}}$, After intersection operator, $M_{\text{private}}=\cap_{i=1}^{n}M^{i}-\cap_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(B), the green-red node is categorized as a privacy-threatening answer because it denotes the existence in the answer subsets.
Union: A union operator’s output will be regarded as private if the answer is the element that belongs to the private answer set of computational subgraphs while not belonging to public answer sets. Given a set of answer sets $M^{1},M^{2},...,M^{n}$, $M_{\text{private}}=\cup_{i=1}^{n}M^{i}_{\text{private}}-\cup_{i=1}^{n}M^{i}_{\text{public}}$. As shown in Figure 3(C), all the red nodes within the answer sets can only be inferred from private attribute links and are classified as privacy-threatening answers.
5. Privacy-Preserving Neural Graph Database
In this section, we present the privacy-preserving neural graph databases for protecting sensitive information in knowledge graphs while preserving high-quality complex query answering performance. We set two optimization goals for P-NGDBs: preserve the accuracy of retrieving non-private answers and obfuscate the privacy-threatening answers.
5.1. Encoding representation
The query encoding methods encode queries to embedding space and compare the similarity with the entities or attributes for answer retrieval, the encoding process can be represented as:
(2) $q=f_{g}(Query),$
where $f_{g}$ is the query encoding methods parameterized by $\theta_{g}$, and $q\in R^{d}$ represents the query embedding for simplicity. Given a query, we iteratively compute the queries based on the sub-query embeddings and logical operators. Assume the encoding process in step $i$, the subquery embedding is $q_{i}$. We can denote the logical operators projection, intersection, and union as:
(3) $q_{i+1}$ $=f_{p}(q_{i},r),\quad r\in\mathcal{R}\cup\mathcal{A},$
(4) $q_{i+1}$ $=f_{i}(q^{1}_{i},...,q^{n}_{i}),$
(5) $q_{i+1}$ $=f_{u}(q^{1}_{i},...,q^{n}_{i}),$
where $f_{p}$, $f_{i}$, and $f_{u}$ denote the parameterized projection, intersection, and union operators, respectively. After query encoding, we compute the score for every candidate based on the query encoding and entity (attribute) embeddings. Finally, we calculate the normalized probability using the Softmax function:
(6) $p(q,v)=e^{s(q,e_{v})}/\sum_{u\in\mathcal{C}}e^{s(q,e_{u})},$
where $s$ is the scoring function, like similarity, distance, etc., and $\mathcal{C}$ is the candidate set.
5.2. Learning Objective
In privacy-preserving neural graph databases, there are two learning objectives: given a query, P-NGDBs should accurately retrieve non-private answers and obfuscate the private answers. For public retrieval, given query embedding $q$, the loss function can be expressed as:
(7) $L_{u}=-\frac{1}{N}\sum_{v\in\mathcal{N}}\log p(q,v),$
where $\mathcal{N}$ is the public answer set for the given query.
While for privacy protection, the learning objective is antithetical. Instead of retrieving correct answers, the objective is to provide obfuscated answers to address privacy concerns. Therefore for private answers, given query embedding $q$, the objective can be expressed as:
(8) $\hat{\theta}_{g}=\arg\min_{\theta_{g}}\sum_{v\in\mathcal{M}_{\text{private}}^{q}}\log p(q,v),$
where $\mathcal{M}_{\text{private}}^{q}$ is the private risk answer set for given query.
As discussed in Section 4.2, the query can be decomposed into sub-queries and logical operators. For intersection and union, the logical operators do not generate new privacy invasive answers, while for projection, the answer can be private if the projection is involvement of private information. Therefore, we can directly optimize the projection operators to safeguard sensitive information. The privacy protection learning objective can be expressed as:
(9) $L_{p}=\frac{1}{|\mathcal{A}_{\text{private}}|}\sum_{r(u,v)\in\mathcal{A}_{\text{private}}}\log p(f_{p}(e_{v},r),u).$
We aim to reach the two objectives simultaneously and the final learning objective function can be expressed as:
(10) $L=L_{u}+\beta L_{p},$
where $\beta$ is the privacy coefficient controlling the protection strength in P-NGDBs, larger $\beta$ denotes stronger protection.
6. Experiments
In this section, to evaluate the privacy leakage problem in NGDBs and the protection ability of our proposed Privacy-preserving Neural Graph Database (P-NGDBs), we construct a benchmark on three real-world datasets and evaluate P-NGDB’s performance based on that.
Figure 4. Eight general query types. Black, blue, and orange arrows denote projection, intersection, and union operators respectively.
6.1. Datasets
We create our benchmarks based on the numerical complex query datasets proposed in (Bai et al., 2023a), which is constructed on FB15K-numerical, DB15K-numerical, and YAGO15K-numerical (Kotnis and García-Durán, 2018), three publicly available knowledge graphs datasets. In each knowledge graph, there are vertices describing entities and attributes and edges describing entity relations, entity attributes, and numerical relations. We first randomly select a set of edges denoting entities’ attributes as privacy information. Then We divide the remaining edges with a ratio of 8:1:1 to construct training, validation, and testing edge sets, respectively. Then the training graph $\mathcal{G}_{train}$, validation graph $\mathcal{G}_{val}$, and testing graph $\mathcal{G}_{test}$ are constructed on training edges, training+validation edges and training+validation+testing edges, respectively. The detailed statistics of the three knowledge graphs are shown in Table 1, #Nodes represents the number of entities and attributes. #Edeges represents the number of relation triples and attribute triples. #Pri. Edges represents the number of attribute triples that are considered private.
Table 1. The statistics of the three knowledge graphs and sampled private attributes.
Graphs Data Split #Nodes #Edges #Pri. Edges
FB15k-N Training 22,964 1,037,480 8,000
Validation 24,021 1,087,296
Testing 27,144 1,144,506
DB15k-N Training 27,639 340,936 6,000
Validation 29,859 381,090
Testing 36,358 452,348
YAGO15k-N Training 30,351 383,772 1,600
Validation 31,543 417,356
Testing 33,610 453,688
Table 2. The statistics of the testing queries’ answers sampled from the three knowledge graphs.
Graphs #Test Ans. 1p 2p 2i 3i pi ip 2u up All
FB15k-N Public 42,952 768,012 138,446 68,677 1,120,065 307,546 1,247,120 1,844,682 5,537,500
Private 524 65,212 1,637 2,576 322,928 30,927 779 239,176 663,099
DB15k-N Public 16,776 764,899 26,836 20,426 1,050,883 105,096 1,296,453 1,959,063 5,240,432
Private 211 25,236 677 763 148,769 6,067 266 146,218 328,207
YAGO15k-N Public 14,880 913,082 38,546 23,066 1,553,590 164,430 1,111,663 1,920,157 5,739,414
Private 136 2,319 499 1742 3,656 1,483 175 40,267 50,277
6.2. Benchmark Construction
Following previous work (Hamilton et al., 2018; Bai et al., 2023a), we evaluate the complex query answering performance on the following eight general query types with abbreviations of $1p,2p,2i,3i,ip,ip,2u$, and $up$, as shown in Figure 4. For each general query type, each edge represents either a projection or a logical operator, and each node represents either a set of entities or numerical values. We use the sampling method proposed in (Bai et al., 2023a) to randomly sample complex queries from knowledge graphs. We random sample training, validation, and testing queries from the formerly constructed graphs respectively. For the training queries, we search for corresponding training answers on the training graph. For the validation queries, we use those queries that have different answers on the validation graph from answers on the training graph to evaluate the generalization ability. For the testing queries, we conduct a graph search on the testing graph with private edges for testing answers and we split the answers into private set and public set according to the definition of privacy risk query answers discussed in Section 4. We conduct a statistical analysis of the number of privacy risk answers for different types of complex queries of three knowledge graphs and the statistics are shown in Table 2.
6.3. Experimental Setup
6.3.1. Baselines
. Our proposed P-NGDBs can be applied to various complex query encoding methods to provide privacy protections. We select three commonly used complex query encoding methods and compare the performance with and without P-NGDB’s protection.
  • GQE (Hamilton et al., 2018): the graph query encoding model encodes a complex query into a vector in embedding space;
  • Q2B (Ren et al., 2020): the graph query encoding model encodes a complex query into a hyper-rectangle embedding space.
  • Q2P (Bai et al., 2022): the graph query encoding model encodes a complex query into an embedding space with multiple vectors.
To the best of our knowledge, there are no privacy protection methods in NGDBs. Therefore, we compare our methods with noise disturbance which is similar to differential privacy (Kasiviswanathan et al., 2013; Qin et al., 2017), a commonly employed technique in database queries that introduces randomness to answers through the addition of noise.
6.3.2. Evaluation Metrics
The evaluation consists of two distinct parts: reasoning performance evaluation and privacy protection evaluation. In performance evaluation, following the previous work (Bai et al., 2023a), we evaluate the generalization capability of models by calculating the rankings of answers that cannot be directly retrieved from an observed knowledge graph. Given a testing query $q$, the training, validation, and public testing answers are denoted as $M_{train}$, $M_{val}$, and $M_{test}$, respectively. We evaluate the quality of retrieved answers using Hit ratio (HR) and Mean reciprocal rank (MRR). HR@K metric evaluates the accuracy of retrieval by measuring the percentage of correct hits among the top K retrieved items. The MRR metric evaluates the performance of a ranking model by computing the average reciprocal rank of the first relevant item in a ranked list of results. The metric can be defined as:
(11) $\text{Metric}(q)=\frac{1}{|M_{test}/M_{val}|}\sum_{v\in M_{test}/M_{val}}m(rank(v)),\$
where $m(r)=\textbf{1}[r\leq K]$ if the metric is HR@K and $m(r)=\frac{1}{r}$ if the metric is MRR. Higher values denote better reasoning performance. In privacy protection evaluation, we compute the rankings metric of privacy-threatening answers as these answers cannot be inferred from the observed graphs, and smaller values denote stronger protection. We train all the models by using the training queries and private attributes and tune hyper-parameters using the validation queries. The evaluation is then finally conducted on the testing queries, including the evaluation of public answers for performance assessment and the evaluation of private answers for privacy assessment. The experiment results are reported on the testing public and private queries respectively.
Table 3. The main experiment results of public answers retrieval and private answers protection. NGDB can provide effective protection while sacrificing acceptable performance.
Dataset Encoding Model Public Private
HR@3 MRR HR@3 MRR
FB15k-N GQE Baseline 21.99 20.26 28.99 27.82
Noise 15.89 14.67 21.54 21.37
P-NGDB 15.92 14.73 10.77 10.21
Q2B Baseline 18.70 16.88 30.28 28.98
Noise 12.34 12.19 20.01 19.71
P-NGDB 12.28 11.18 10.17 9.38
Q2P Baseline 26.45 24.48 29.08 31.85
Noise 20.13 19.77 22.35 23.17
P-NGDB 19.48 18.19 14.15 14.93
DB15k-N GQE Baseline 24.16 22.37 39.26 37.25
Noise 18.01 16.35 28.59 28.37
P-NGDB 17.58 16.29 10.52 10.79
Q2B Baseline 15.94 14.98 42.19 39.78
Noise 10.76 10.28 26.49 25.93
P-NGDB 10.19 9.49 8.92 7.99
Q2P Baseline 25.72 24.12 46.18 43.48
Noise 19.89 19.32 33.56 33.17
P-NGDB 20.26 19.00 19.38 18.45
YAGO15k-N GQE Baseline 26.06 24.37 43.55 40.81
Noise 20.32 20.27 38.52 38.29
P-NGDB 19.58 19.82 7.56 7.33
Q2B Baseline 23.39 22.53 42.73 40.55
Noise 16.85 15.37 28.23 28.54
P-NGDB 17.07 16.03 6.26 5.79
Q2P Baseline 29.41 27.87 42.56 45.79
Noise 22.85 21.21 34.26 33.68
P-NGDB 23.27 22.59 7.34 7.17
Table 4. The mean reciprocal ranking (MRR) variation for P-NGDBs on different types of queries.
Dataset Enc. Test. 1p 2p 2i 3i pi ip 2u up
FB15k-N GQE Public 10.51(70.4%) 3.89(55.3%) 23.49(89.1%) 44.14(92.2%) 14.57(83.4%) 6.43(73.5%) 4.15(59.2%) 2.95(70.9%)
Private 0.18(4.6%) 1.80(13.6%) 11.77(45.7%) 32.71(56.2%) 14.07(46.6%) 6.37(36.7%) 0.70(6.1%) 1.03(13.6%)
Q2B Public 12.88(69.8%) 3.27(53.6%) 20.13(85.3%) 36.52(83.4%) 12.18(69.7%) 4.57(59.1%) 5.41(57.8%) 3.04(62.5%)
Private 0.41(11.6%) 3.02(23.4%) 9.45(31.5%) 27.42(50.3%) 15.21(47.5%) 9.49(47.1%) 0.23(5.0%) 1.48(19.1%)
Q2P Public 16.36(73.5%) 6.22(59.8%) 23.34(82.5%) 43.29(93.6%) 15.26(77.3%) 7.44(63.8%) 6.79(60.8%) 3.64(58.2%)
Private 0.31(11.3%) 3.24(31.3%) 15.82(46.7%) 39.49(46.7%) 18.12(52.7%) 9.60(45.1%) 0.38(8.1%) 2.55(23.3%)
DB15k-N GQE Public 2.35(58.1%) 0.98(56.8%) 22.30(83.1%) 49.54(85.4%) 10.06(67.0%) 3.48(53.0%) 0.73(51.2%) 0.94(57.2%)
Private 0.22(9.3%) 1.76(17.4%) 12.33(17.7%) 31.88(44.9%) 23.26(57.9%) 4.83(47.9%) 0.76(12.1%) 0.94(19.2%)
Q2B Public 2.38(53.0%) 1.54(65.8%) 16.16(83.5%) 28.44(85.1%) 9.23(63.9%) 2.68(52.3%) 0.81(57.2%) 0.98(55.4%)
Private 0.18(7.6%) 2.72(26.8%) 9.34(13.4%) 24.89(34.7%) 19.10(53.3%) 6.06(53.0%) 0.74(12.5%) 0.67(16.0%)
Q2P Public 4.75(60.3%) 2.59(61.8%) 24.67(80.2%) 52.48(94.8%) 13.41(70.2%) 4.62(48.1%) 2.20(54.3%) 1.58(46.8%)
Private 0.23(8.7%) 2.92(28.4%) 28.37(39.5%) 35.58(47.0%) 24.55(57.4%) 9.05(57.6%) 1.62(23.9%) 1.87(22.5%)
YAGO15k-N GQE Public 5.28(86.5%) 1.78(57.9%) 24.35(66.9%) 43.2(65.0%) 13.98(70.3%) 4.32(60.7%) 3.56(89.2%) 1.03(66.0%)
Private 0.12(6.3%) 3.21(15.0%) 12.56(37.2%) 27.65(35.6%) 14.09(45.2%) 12.87(53.7%) 1.08(7.9%) 1.20(9.3%)
Q2B Public 6.06(74.3%) 1.98(51.5%) 24.03(72.2%) 44.52(81.7%) 13.22(63.3%) 4.33(52.4%) 3.84(61.9%) 1.11(56.1%)
Private 0.11(5.6%) 3.45(16.6%) 16.99(50.4%) 20.62(32.4%) 13.51(39.1%) 14.42(48.3%) 1.32(10.1%) 1.14(12.3%)
Q2P Public 9.96(83.7%) 3.78(67.7%) 30.96(77.1%) 37.87(68.9%) 15.23(64.3%) 6.59(64.9%) 9.63(71.4%) 2.18(66.1%)
Private 0.12(9.0%) 3.17(18.1%) 14.29(35.0%) 34.29(42.6%) 16.59(45.2%) 13.95(47.2%) 1.96(9.9%) 2.68(13.6%)
Avg. Change Public $\downarrow$ 30.0% $\downarrow$ 41.1% $\downarrow$ 20.0% $\downarrow$ 16.7% $\downarrow$ 30.1% $\downarrow$ 41.4% $\downarrow$ 37.4% $\downarrow$ 40.1%
Private $\downarrow$ 91.8% $\downarrow$ 78.8% $\downarrow$ 59.9% $\downarrow$ 56.7% $\downarrow$ 50.6% $\downarrow$ 51.5% $\downarrow$ 67.1% $\downarrow$ 83.5%
6.3.3. Parameter Setting.
We first tune hyper-parameters on the validation queries for the base query encoding methods. We use the same parameters for NGDBs for a fair comparison. We tune the privacy penalty coefficient $\beta$ for three datasets respectively for utility privacy tradeoff for convenient illustration. For noise disturbance privacy protection, we adjust the noise strength to make the results comparable to NGDBs’ results (similar performance on public query answers).
6.4. Performance Evaluation
We apply our P-NGDB on three different query encoding methods GQE (Hamilton et al., 2018), Q2B (Ren et al., 2020), and Q2P (Bai et al., 2022) and compare the query answering performance with and without P-NGDB’s protection. Experiment results are summarized in Table 3 and Table 4. Table 3 reports the averaged results in HR@3 and MRR. A higher value on public answer sets denotes better reasoning ability, while a smaller value on private answer sets denotes stronger protection. The experiment results show that without protection, the base encoding methods all suffer from privacy leakage problems if there is sensitive information existing in the knowledge graph. Our proposed P-NGDB can effectively protect private information with a slight loss of complex query answering ability. For example, in FB15K-N, GQE retrieves the public answers with 21.99 HR@3 and the private answers with 28.99 HR@3 without privacy protection. With P-NGDB’s protection, GQE retrieves public answers with 15.92 HR@3 and private answers with 10.77 HR@3. The private answers are protected with a 62.9% decrease in HR@3 sacrificing a 27.4% decrease in HR@3 for public answers. Compared to the noise disturbance protection, our method can accurately protect sensitive information, so that the loss of performance is much lower than the noise disturbance.
In Table 4, we show P-NGDBs’ MRR results on various query types and the percentages within parentheses represent the performance of the P-NGDB protected model relative to the performance of the encoding model without protection. The evaluation is conducted on public and private answers respectively to assess privacy protection. From the comparison between MRR variation on public and private answers, we can know that P-NGDB can provide protection on all the query types. Besides, comparing the average change on these query types, we can know from the results that the P-NGDB have different protection ability on various query types. For projection and union, the MRR of protected models on private answers is largely degraded, denoting a stronger protection. While for the intersection operator, the preservation provided by P-NGDBs is as effective as protections for other operators.
Figure 5. The evaluation results of GQE with various privacy coefficients $\beta$ on FB15K-N.
6.5. Sensitivity Study
The privacy protection needs can be different in various knowledge graphs. We evaluate the impact of the privacy penalty coefficient $\beta$, A larger $\beta$ can provide stronger privacy protection. We select $\beta$ from $\{0.01,0.05,0.1,0.5,1\}$, and evaluate the retrieval accuracy of public answers and private answers, respectively. We evaluate GQE’s retrieval performance on the FB15K-N dataset and depict the MRR’s change under each $\beta$ compared to the unprotected GQE model. The results are shown in Figure 5. We can know that the penalty can effectively control the level of privacy protection, to meet higher privacy needs, P-NGDBs have to undertake much utility loss. Besides, the flexible adjustment of the privacy coefficient can make P-NGDBs adapt to various scenarios with different privacy requirements.
7. Conclusion
In this work, we proposed the privacy problem in neural graph databases and showed that sensitive information can be inferred by attackers with specified queries. We denote that some query answers can leak privacy information named privacy risk answer. To systematically evaluate the problem, we constructed a benchmark dataset based on FB15k-N, YAGO15k-N, and DB15k-N. Finally, we proposed a new framework that can protect the privacy of the knowledge graph from attackers’ malicious queries. Experimental results on the benchmark shows that the proposed model (P-NGDB) can effectively protect privacy while sacrificing slight reasoning performance. In the future, we will take logical operators into consideration and make improvements in NGDB privacy protection.